OpenStudy (anonymous):

how many real-number solutions does this equation have? -8x^2-8x-2=0 a. one solution b. two solutions c. no solutions d. infinetly many solutions

OpenStudy (solomonzelman):

To answer the question `how many real-number solutions does the equation have` I would solve it. \(\normalsize\color{blue}{ -8x^2-8x-2=0 }\) factor the left side out of `-2` \(\normalsize\color{blue}{ -2(4x^2+4x+1)=0 }\) divide both sides by `-2` \(\normalsize\color{blue}{ 4x^2+4x+1=0 }\) \(\normalsize\color{blue}{ (2x+1)^2=0 }\) square root both sides, \(\normalsize\color{blue}{ 2x+1=0 }\)

OpenStudy (solomonzelman):

Let me know if you need more help solving the equation, and determining the number of REAL solutions.

OpenStudy (anonymous):

no thanks! thank you so much that helped alot !!!!!!!!! :)))) @SolomonZelman

OpenStudy (solomonzelman):

Anytime, so how many (real) solutions are there ?